# EEG-MACS: Manifold Attention and Confidence Stratification for EEG-based Cross-Center Brain Disease Diagnosis under Unreliable Annotations

### v1

@@ 3.3.1. Multi-view Contrastive Loss
3) $\mathcal{L}^{St}$: This component incorporates interpolation supervised contrastive learning techniques, as exploited in [40, 22], and shares the same hyperparameter $\lambda$ with Eq.(3). It effectively combines $\mathcal{L}^{St}_{i\in\Psi^{*}}$ and $\mathcal{L}^{St}_{i\in\Psi^{+}}$ from trusted pairs $\Psi=\Psi^{*}\cup\Psi^{+}$. It can be expressed as $\mathcal{L}_{i}^{St}=\lambda\mathcal{L}_{i}\left(z_{i},y^{*}\right)+(1-\lambda)\mathcal{L}_{i}\left(z_{i},y^{+}\right)$.
@@ 3.3.2. Discriminative Loss
Frequency[7] 58.64(2.06) 66.03(1.47) 74.07(0.07) 71.90(0.02) 66.36 71.51
Entropy[8] 63.07(1.46) 69.89(1.50) 61.11(0.21) 57.68(2.29) 62.09 72.37
Hjorth[6] 67.05(0.80) 68.37(0.75) 64.81(0.27) 63.78(0.49) 65.93 64.65
Functional Connectivity-Corr[41] 65.34(1.64) 67.56(2.58) 66.67(0.21) 61.39(0.78) 66.01 63.99
Functional Connectivity-SPD [36] 64.43(0.77) 68.18(0.99) 79.63(0.27) 79.46(0.17) 72.03 74.80
@@
[10] [Ju et al., TNNLS, 2022] Tensor-CSPNet 80.78(7.93) 77.70(11.63) 75.92(6.41) 73.20(8.16) 78.35(3.44)
[11] [Pan et al., NeurIPS, 2022] MAtt 81.97(3.96) 79.56(8.29) 79.63(11.56) 83.07(7.56) 80.80(1.65)
[3] [Wang et al., NeurIPS, 2023] COMET 73.25(26.02) 70.09(24.82) 75.47(9.75) 74.92(14.17) 74.36(1.57)
@@
[20][Ma et al., IJCAI, 2023] CTW 75.40(10.28) 76.75(9.80) 75.93(3.21) 78.73(2.20) 75.67(0.37)
[21][Xiao et al., IJCAI, 2023] Promix 55.14(18.94) 56.17(11.49) 79.63(8.49) 77.34(10.97) 67.39(17.32)
[20] [Ma et al., IJCAI, 2023]
[21] [Xiao et al., IJCAI, 2023]
@@
[22] [Li et al., CVPR, 2022] Sel-CL 63.29(22.32) 54.82(19.42) 66.67(5.56) 65.42(4.49) 64.98(2.39)
[22] [Li et al., CVPR, 2022] Sel-CL+ 74.16(7.73) 73.41(11.98) 57.41(21.03) 43.06(37.35) 65.79(11.84)
[21][Xiao et al., IJCAI, 2023] Encoder for Promix 85.38(4.41) 85.56(4.90) 83.33(5.56) 81.05(8.49) 84.36(1.45)
[22] [Li et al., CVPR, 2022] Encoder for Sel-CL 72.08(12.76) 70.05(9.37) 83.33(5.56) 81.51(9.04) 77.71(7.95)
[22] [Li et al., CVPR, 2022] Encoder for Sel-CL+ 85.43(4.16) 84.18(7.01) 83.33(5.56) 81.05(8.49) 84.38(1.48)
@@ 4.1. Experimental Datasets
The PD data includes two sets. The first set, from the University of New Mexico(NMU), comprises data from 27 patients and 27 controls [7]. The second set, from the University of Iowa(IU), includes data from 14 patients and 14 controls [42]. Both datasets were acquired using 64-channel Ag/AgCl electrodes with the Brain Vision system at a sampling rate of 500 Hz.
@@ 4.3. Feature Engineering Study
Feature engineering insights have led to the development of the MACS Encoder. EEG frequency markers [7]are inductive for PD but found less effective for MCI. EEG complexity measured by the entropy[8] and Hjorth parameter[6] enhanced MCI detection but was less effective than frequency domain features in PD cases. We compared the functional connectivity characterized by Pearson correlation (Corr) in Euclidean space [41] and SPD in manifold space [36] and found the advantage of manifold representations. This evidence has informed the integration of a specialized block in the Encoder to process underlying temporal-spectral-spatial features and analyze synchronization characteristics through manifold geometry.
@@ 4.4. Comparison Study
We benchmarked MACS against SOTA methods, evaluating its performance in scenarios with unreliable annotations and independently assessing the Encoder in scenarios with clean annotations (see Table 2). For the clean data learning comparison, we employed advanced EEG encoders that target manifold geometry[10], integrate attention mechanisms[11], and utilize contrastive learning[3]. Given the absence of a dedicated SOTA model for EEG data with unreliable annotations, we compared MACS against leading methods in the time series[20] and natural image domain[21, 22]. Specifically, we replaced their encoders with our Encoder in MACS to validate the efficacy of our combined manifold-Euclidean representation learning strategy for EEG signals. Overall, MACS’s weakly supervised learning mechanism, combined with an effective Encoder, enables superior comprehensive performance across two disease types. We unveil MACS’s learning process through t-SNE mapping in Figure 5(see Appendix for the PD case). Observations indicate that MACS gradually selects the correct trust samples, enhancing representation learning and forming class-specific clusters.
@@ 4.5. Hyperparameter Tuning Study
The multi-view contrastive loss is crucial for the MACS framework. Effective contrastive learning heavily depends on the adequacy of positive and negative pairs; therefore, inspired by [40], we proposed maintaining a memory of previous steps to circumvent the constraints imposed by batch size, enabling more extensive use of available small-scale data. For example, we summarize the research findings for MCI in Table 3, suggesting that memory length contributes to learning efficiency, but its relationship is not monotonic. Please note that the methods involving contrastive learning are discussed in Table 2, where we also determined their optimal performance to ensure a fair comparison.
@@
Temporal Scale Temporal scale refers to the time interval for constructing dynamic functional networks in the Encoder. The findings indicate that a second-level temporal scale is crucial for capturing patterns. For example, results from the MCI dataset are included in Table 3. More effective EEG markers that appear at larger temporal scales may be attributed to the ’slowing’ phenomenon in brain activity, which is more prominently manifested in patients with neurodegenerative disorders [43, 44].
@@ 4.7. Transferability Evaluation
From cross-validation tests on PD(NMU) and MCI datasets, we derived N-fold MACS models for PD ($\Theta_{PD}$) and MCI ($\Theta_{MCI}$). We assessed MACS’s transferability by testing $\{\theta_{PD}^{i}\}_{i=0}^{2}$ on a cross-center PD corpus (from IU), and fine-tuning $\{\theta_{MCI}^{i}\}_{i=0}^{3}$ on a cross-center AD corpus. To address the spatial resolution gap between MCI and AD data, we employed EEG’s interpolation technique [45]. The averaged evaluation results from $\Theta_{PD}$, halted at both optimal and final epochs (Figure 6(a)), and successful fine-tuning on AD with few labels (Figure 6(b)), underscore MACS’s notable transferability in recognizing neurodegenerative disorders via EEG.

### v2

@@ 3.3.1. Multi-view Contrastive Loss
3) $\mathcal{L}^{St}$: This component incorporates interpolation supervised contrastive learning techniques, as exploited in (Ortego et al., 2021; Li et al., 2022), and shares the same hyperparameter $\lambda$ with Eq.(3). It effectively combines $\mathcal{L}^{St}_{i\in\Psi^{*}}$ and $\mathcal{L}^{St}_{i\in\Psi^{+}}$ from trusted pairs $\Psi=\Psi^{*}\cup\Psi^{+}$. It can be expressed as $\mathcal{L}_{i}^{St}=\lambda\mathcal{L}_{i}\left(z_{i},y^{*}\right)+(1-\lambda)\mathcal{L}_{i}\left(z_{i},y^{+}\right)$.
@@ 3.3.2. Discriminative Loss
Frequency(Cavanagh et al., 2018) 58.64(2.06) 66.03(1.47) 74.07(0.07) 71.90(0.02) 66.36 71.51
Entropy(Suuronen et al., 2023) 63.07(1.46) 69.89(1.50) 61.11(0.21) 57.68(2.29) 62.09 72.37
Hjorth(Coelho et al., 2023) 67.05(0.80) 68.37(0.75) 64.81(0.27) 63.78(0.49) 65.93 64.65
Functional Connectivity-Corr(Lopez et al., 2023) 65.34(1.64) 67.56(2.58) 66.67(0.21) 61.39(0.78) 66.01 63.99
Functional Connectivity-SPD (Bonet et al., 2023) 64.43(0.77) 68.18(0.99) 79.63(0.27) 79.46(0.17) 72.03 74.80
@@
(Ju and Guan, 2023) [Ju et al., TNNLS, 2022] Tensor-CSPNet 80.78(7.93) 77.70(11.63) 75.92(6.41) 73.20(8.16) 78.35(3.44)
(Pan et al., 2022) [Pan et al., NeurIPS, 2022] MAtt 81.97(3.96) 79.56(8.29) 79.63(11.56) 83.07(7.56) 80.80(1.65)
(Wang et al., 2023a) [Wang et al., NeurIPS, 2023] COMET 73.25(26.02) 70.09(24.82) 75.47(9.75) 74.92(14.17) 74.36(1.57)
@@
(Ma et al., 2023)[Ma et al., IJCAI, 2023] CTW 75.40(10.28) 76.75(9.80) 75.93(3.21) 78.73(2.20) 75.67(0.37)
(Xiao et al., 2023)[Xiao et al., IJCAI, 2023] Promix 55.14(18.94) 56.17(11.49) 79.63(8.49) 77.34(10.97) 67.39(17.32)
(Ma et al., 2023) [Ma et al., IJCAI, 2023]
(Xiao et al., 2023) [Xiao et al., IJCAI, 2023]
@@
(Li et al., 2022) [Li et al., CVPR, 2022] Sel-CL 63.29(22.32) 54.82(19.42) 66.67(5.56) 65.42(4.49) 64.98(2.39)
(Li et al., 2022) [Li et al., CVPR, 2022] Sel-CL+ 74.16(7.73) 73.41(11.98) 57.41(21.03) 43.06(37.35) 65.79(11.84)
(Xiao et al., 2023)[Xiao et al., IJCAI, 2023] Encoder for Promix 85.38(4.41) 85.56(4.90) 83.33(5.56) 81.05(8.49) 84.36(1.45)
(Li et al., 2022) [Li et al., CVPR, 2022] Encoder for Sel-CL 72.08(12.76) 70.05(9.37) 83.33(5.56) 81.51(9.04) 77.71(7.95)
(Li et al., 2022) [Li et al., CVPR, 2022] Encoder for Sel-CL+ 85.43(4.16) 84.18(7.01) 83.33(5.56) 81.05(8.49) 84.38(1.48)
@@ 4.1. Experimental Datasets
The PD data includes two sets. The first set, from the University of New Mexico(NMU), comprises data from 27 patients and 27 controls (Cavanagh et al., 2018). The second set, from the University of Iowa(IU), includes data from 14 patients and 14 controls (Anjum et al., 2020). Both datasets were acquired using 64-channel Ag/AgCl electrodes with the Brain Vision system at a sampling rate of 500 Hz.
@@ 4.3. Feature Engineering Study
Feature engineering insights have led to the development of the MACS Encoder. EEG frequency markers (Cavanagh et al., 2018)are inductive for PD but found less effective for MCI. EEG complexity measured by the entropy(Suuronen et al., 2023) and Hjorth parameter(Coelho et al., 2023) enhanced MCI detection but was less effective than frequency domain features in PD cases. We compared the functional connectivity characterized by Pearson correlation (Corr) in Euclidean space (Lopez et al., 2023) and SPD in manifold space (Bonet et al., 2023) and found the advantage of manifold representations. This evidence has informed the integration of a specialized block in the Encoder to process underlying temporal-spectral-spatial features and analyze synchronization characteristics through manifold geometry.
@@ 4.4. Comparison Study
We benchmarked MACS against SOTA methods, evaluating its performance in scenarios with unreliable annotations and independently assessing the Encoder in scenarios with clean annotations (see Table 2). For the clean data learning comparison, we employed advanced EEG encoders that target manifold geometry(Ju and Guan, 2023), integrate attention mechanisms(Pan et al., 2022), and utilize contrastive learning(Wang et al., 2023a). Given the absence of a dedicated SOTA model for EEG data with unreliable annotations, we compared MACS against leading methods in the time series(Ma et al., 2023) and natural image domain(Xiao et al., 2023; Li et al., 2022). Specifically, we replaced their encoders with our Encoder in MACS to validate the efficacy of our combined manifold-Euclidean representation learning strategy for EEG signals. Overall, MACS’s weakly supervised learning mechanism, combined with an effective Encoder, enables superior comprehensive performance across two disease types. We unveil MACS’s learning process through t-SNE mapping in Figure 5 (see Appendix for the PD case). Observations indicate that MACS gradually selects the correct trust samples, enhancing representation learning and forming class-specific clusters.
@@ 4.5. Hyperparameter Tuning Study
The multi-view contrastive loss is crucial for the MACS framework. Effective contrastive learning heavily depends on the adequacy of positive and negative pairs; therefore, inspired by (Ortego et al., 2021), we proposed maintaining a memory of previous steps to circumvent the constraints imposed by batch size, enabling more extensive use of available small-scale data. For example, we summarize the research findings for MCI in Table 3, suggesting that memory length contributes to learning efficiency, but its relationship is not monotonic. Please note that the methods involving contrastive learning are discussed in Table 2, where we also determined their optimal performance to ensure a fair comparison.
@@
Temporal Scale Temporal scale refers to the time interval for constructing dynamic functional networks in the Encoder. The findings indicate that a second-level temporal scale is crucial for capturing patterns. For example, results from the MCI dataset are included in Table 3. More effective EEG markers that appear at larger temporal scales may be attributed to the ’slowing’ phenomenon in brain activity, which is more prominently manifested in patients with neurodegenerative disorders (Song et al., 2018; Jones et al., 2022).
@@ 4.7. Transferability Evaluation
From cross-validation tests on PD(NMU) and MCI datasets, we derived N-fold MACS models for PD ($\Theta_{PD}$) and MCI ($\Theta_{MCI}$). We assessed MACS’s transferability by testing $\{\theta_{PD}^{i}\}_{i=0}^{2}$ on a cross-center PD corpus (from IU), and fine-tuning $\{\theta_{MCI}^{i}\}_{i=0}^{3}$ on a cross-center AD corpus. To address the spatial resolution gap between MCI and AD data, we employed EEG’s interpolation technique (Gil Ávila et al., 2023). The averaged evaluation results from $\Theta_{PD}$, halted at both optimal and final epochs (Figure 6(a)), and successful fine-tuning on AD with few labels (Figure 6(b)), underscore MACS’s notable transferability in recognizing neurodegenerative disorders via EEG.
@@ 5. Conclusion
Acknowledgements.
This work is supported by the National Natural Science Foundation of China (Grant No. 62306089), the Shenzhen Science and Technology Program (Grant No. RCBS20231211090800003), and the China Postdoctoral Science Foundation (Grant Nos. 2023M730873 and GZB20230960).